What is the perimeter of a rectangle if the area is 36 m square?
Found 2 solutions by JimboP1977, checkley77:
Answer by JimboP1977(311)

Impossible to say for certain. The dimensions could be 9 by 4 for example which gives a perimeter of 26 metres. Or it could be 3 by 12, which gives a perimeter of 30 metres...

Answer by checkley77(12844)

LW=36
THERE ARE MANY SOLUTIONS.
A FEW FOLLOW:
L*W/////2L+2W
6*6///2*6+2*6=12=12=24 ANS.
9*4///2*9+2*4=18+8=26 ANS.
12*3//2*12+2*3=24+6=30 ANS.
18*2//2*18+2*2=36+4=40 ANS.
36*1//2*36+2*1=72+2=74 ANS.